# Gravitational mode mixing around black holes in scalar-tensor theories with parity-violating terms

### v1

@@ Abstract
We investigate black holes and gravitational perturbations when both the scalar Gauss-Bonnet and dynamical Chern-Simons gravity sectors coexist in addition to the Einstein-Hilbert term, and both sectors are coupled to a single canonically normalized scalar field. The presence of the scalar Gauss-Bonnet gravity sector allows the scalar field to possess a non-vanishing background solution, resulting in additional couplings between odd and even-type gravitational perturbations arising from the dynamical Chern-Simons gravity sector. We illustrate the impact of these even-odd gravitational couplings in gravitational perturbations around a static spherically symmetric black hole. Although the couplings between the odd and even-type gravitational perturbations are known to appear in purely tensorial gravity theories with higher curvature corrections, we demonstrate it in scalar-tensor theories.
@@ I Introduction
General Relativity (GR) faces several challenges of the late-time acceleration Riess et al. (1998); Perlmutter et al. (1999), the formation of black hole singularities Penrose (1965); Hawking and Penrose (1970); Hawking (1976), and the quantization of gravity at both low and high energy scales. These strongly suggest that the theory of gravity should be modified from general relativity. In most cases, the modification of gravity involves adding extra degrees of freedom to the metric. In fact, the simplest modification is to introduce a non-minimal coupling of the metric to a scalar field, i.e., a scalar-tensor theory. As a typical example, we focus both on the scalar Gauss-Bonnet (sGB) gravity Metsaev and Tseytlin (1987); Kanti et al. (1996); Green and Schwarz (1984); Antoniadis et al. (1992) and the dynamical Chern-Simons (dCS) gravity Lue et al. (1999); Jackiw and Pi (2003); Alexander and Yunes (2009). Both theories of gravity are known as low-energy theories of string theory Mignemi and Stewart (1993); Antoniadis et al. (1994); Mignemi (1995); Kanti et al. (1996). The sGB gravity sector breaks the four-dimensional topological invariance of the Gauss-Bonnet term by coupling the dynamical scalar field, while the dCS gravity has a parity-violating coupling between the scalar field and the Pontryagin term. [^1] We note that while the sGB gravity is a healthy classical theory included in the Horndeski theory which is the most general theory with the second-order equations of motion Kobayashi et al. (2011), the dCS gravity should be regarded as a low-energy effective field theory because of the presence of ghost modes around a static spherically symmetric black hole Motohashi and Suyama (2012) and any spacetime Crisostomi et al. (2018). The black hole solutions and QNMs have been studied in the sGB gravity Yunes and Stein (2011); Blázquez-Salcedo et al. (2016, 2017, 2019); Konoplya et al. (2019); Zinhailo (2019); Churilova and Stuchlik (2020); Blázquez-Salcedo et al. (2020); Pierini and Gualtieri (2021, 2022) and the dCS gravity Yunes and Sopuerta (2008); Cardoso and Gualtieri (2009); Molina et al. (2010); Yunes and Stein (2011); Srivastava et al. (2021); Wagle et al. (2022); Blázquez-Salcedo et al. (2023), respectively, where the deviations from the GR case can be expressed as a series of small couplings. Recently, in the context of the effective field theory (EFT), corrections to black hole solutions and QNMs with small EFT couplings have been studied Cardoso et al. (2018); Cano and Ruipérez (2019); de Rham et al. (2020); Nomura and Yoshida (2022); Cano et al. (2022a); Silva et al. (2023); Cano et al. (2023a); Cayuso et al. (2023); Cano et al. (2023b); Melville (2024).
@@
In this paper, we study black holes and gravitational perturbations when both the sGB and dCS gravity sectors coexist, and both sectors are coupled to a single scalar field. The presence of the sGB gravity sector allows the scalar field to have a background value, leading to additional new couplings between the odd and even-type gravitational perturbations appearing through the dCS gravity sector. We illustrate the impact of the even-odd gravitational couplings in the perturbations around the static spherically symmetric black hole solutions. While the couplings between the odd and even-type gravitational perturbations are known to appear in purely tensorial gravity theories with higher curvature corrections Cardoso et al. (2018); de Rham et al. (2020); Cano et al. (2022b, 2023a), we demonstrate it in scalar-tensor theories.
@@ III Gravitational Perturbations
In this section, we study gravitational perturbations around the static spherically symmetric background solutions $\bar{g}_{\mu\nu}$ and $\bar{\phi}$ in Eqs. (4)-(5) with Eqs. (7)-(14). The perturbed scalar field is given by
@@
where $h_{0},~{}h_{1},H_{0},~{}H_{1},~{}H_{2}$ and $K$ are functions of $r$. The odd (--) mode components $h_{0},~{}h_{1}$ and the even ($+$) mode components $H_{0},~{}H_{1},~{}H_{2},K$ are not independent, but related through the equations of motion. In our setup, the metric perturbations have two helicity-2 modes.
@@ III.1 Master equations
The explicit forms of ${\bm{\Psi}}$, $\Delta{\bm{c}}^{2}$ and ${\bm{V}}_{\rm correction}$ are written in App A. Since $\Delta{\bm{c}}^{2}$ vanishes at the horizon, the propagation speed of each mode coincides with the speed of light. This implies that information inside the horizon cannot propagate outside of the horizon. The non-diagonal components of the potential matrix in Eq. (46) represent the mode mixing. We note that the scalar-odd and scalar-even coupling terms at ${\cal O}(\epsilon^{2})$ in Eq. (46) vanish because the next leading terms appear at ${\cal O}(\epsilon^{3})$. Considering both sGB and dCS couplings, new terms, $V^{-+}_{\rm 2nd}$, can appear in the master equations. In particular, the even-odd gravitational couplings have been known to be a generic feature of purely tensorial gravity theories with higher curvature corrections Cardoso et al. (2018); de Rham et al. (2020); Cano et al. (2022b, 2023a). We find that this is characteristic not only in purely tensorial gravity theories but also when both the sGB term and the dCS term are present. In the context of cosmology, it has been known that the even-odd gravitational couplings can appear in a similar system where both sGB and dCS gravity sectors coexist Satoh et al. (2008).
@@ IV Quasinormal Modes
QNMs are the characteristic oscillation modes of perturbations around black holes. The quasinormal frequencies are defined by the solutions of $\omega$ in the master equations imposing the purely ingoing condition at the horizon and the purely outgoing condition at infinity. The Leaver method Leaver (1985) is used to calculate the quasinormal frequencies. This calculation method is briefly presented for both decoupled and coupled systems. In the decoupled system, the master equation is described by a single master variable $\Psi$
@@
where ${\rm d}r/{\rm d}r_{\ast}=f$. The wave function $\Psi$ of this form satisfies the QNM boundary condition at the horizon thanks to the prefactor $e^{-i\omega r_{\ast}}f^{-2i\omega r_{\rm H}}$. Substituting this ansatz into the master equation (49) yields the recursion relation for the coefficients $a_{j}$. For a given $a_{0}$, the higher order coefficients $a_{j}~{}(j=1,~{}2,~{}\cdots)$ can be determined by the recursion relation666 Since the boundary condition at the horizon has already been imposed, the first term $a_{0}$ determines the solution of the master equation. in the form $a_{j}=G_{j}(\omega)a_{0}$, where $G_{j}(\omega)$ is a function in $j$ and $\omega$. If the wave function (48) satisfies the QNM boundary condition at infinity $f=1$, then the sum of the series $\sum^{\infty}_{j=0}a_{j}$ should converge. Such a solution can be derived approximately by imposing $a_{j}=0$ for sufficiently large $j$. Thus, we need to solve $G_{j}(\omega)=0$ for sufficiently large $j$, and the quasinormal frequencies are given by those solutions Leaver (1985). In GR, for $\ell=2$ and the fundamental modes, the quasinormal frequency for the scalar field is $(0.96728774-0.19359239i)/(2M)$ Berti et al. (2009) where $M$ is the black hole mass, and for the odd and even gravitational modes, it is $(0.74734337-0.17792463i)/(2M)$ Chandrasekhar and Detweiler (1975); Berti et al. (2009). In the case of non-GR gravity theories, if the master equation is expressed as a perturbation from the GR case, the quasinormal frequencies can be obtained as small deviations from the GR values Cardoso et al. (2018, 2019); McManus et al. (2019); Cano and Ruipérez (2019); de Rham et al. (2020); Kimura (2020); Nomura and Yoshida (2022); Cano et al. (2022a); Franchini and Völkel (2023); Cano et al. (2023a); Cayuso et al. (2023); Cano et al. (2023b); Hatsuda and Kimura (2024).
In a coupled system, the master equations have non-diagonal parts of the potential matrix777 Changing the master variable, the master equations (27) can be written in the form of Eq. (49) (see App. A in Cardoso et al. (2019)).
@@
where $\Omega^{i}_{\rm Sch}~{}(i={\rm s,~{}grav.})$ are the dimensionless Schwarzschild quasinormal frequencies normalized by the Schwarzschild radius for the scalar field, the odd and even gravitational modes.888 Practically, we first calculate $\delta\bar{\omega}_{i}$ with $\omega=\Omega_{\rm Sch}^{i}/r_{\rm H}+\epsilon^{2}\delta\bar{\omega}_{i}$, and then we can obtain $\delta\omega_{i}$ in Eq. (52) as $\delta\omega_{i}=\delta\bar{\omega}_{i}-\delta R_{\rm H}\delta\Omega_{\rm Sch}^{i}/(2M)$, where $r_{\rm H}=2M(1+\epsilon^{2}\delta R_{\rm H})$. For the fundamental modes, the values of $\Omega^{i}_{\rm Sch}$ are given by $\Omega^{\rm s}_{\rm Sch}=0.96728774-0.19359239i$ and $\Omega^{\rm grav}_{\rm Sch}=0.74734337-0.17792463i$. Substituting these ansatzes into Eq. (51) in our system and expanding them with respect to $\epsilon$, we obtain the equations determining $\delta\omega_{i}$.
@@
where $F_{0},F_{2,0},F_{2,1}$ are functions of $\Omega^{\rm s}_{\rm Sch}$. The lowest order value $\Omega^{\rm s}_{\rm Sch}$ is determined by the equation $F_{0}=0$. The second order ${\cal O}(\epsilon^{2})$ expression of the equation (53) determines the value of $\delta\omega_{\rm s}$. We note that the term $F_{2,1}$ is the derivative of $F_{0}$ with respect to the frequency, and the term $F_{2,0}$ is affected by the correction terms $V^{\rm s}_{\rm 2nd}$ and $V^{\rm s\pm}_{\rm 1st}$ in Eq. (46). The QNM correction $\delta\omega_{\rm s}$ is given by
@@
The corrections to the lowest order values of the odd and even gravitational modes are determined by the self-couplings of those modes $V^{\rm-}_{\rm 2nd}$ and $V^{\rm+}_{\rm 2nd}$, and the even-odd gravitation coupling $V^{\rm-+}_{\rm 2nd}$. Eq. (51) is given by
@@
The terms $\tilde{F}_{0},\tilde{F}_{2,0},\tilde{F}_{2,1},\tilde{F}_{4,0},\tilde{F}_{4,1},\tilde{F}_{4,2}$ are functions of $\Omega^{\rm grav}_{\rm Sch}$. We solve the above equation order by order in $\epsilon$. The lowest-order quasinormal frequency of the gravitational modes $\Omega^{\rm grav}_{\rm Sch}$ is determined by $\tilde{F}_{0}=0$. Since the spectra of even and odd gravitational modes are degenerate at the lowest order, the terms $\tilde{F}_{2,0}$ and $F_{2,1}$, which are proportional to the first derivative of $\tilde{F}_{0}$, also vanish. Thus, the QNM correction $\delta\omega_{\rm grav}$ is determined from Eq. (55) with ${\cal O}(\epsilon^{4})$. We note that the term $\tilde{F}_{4,2}$ is the second derivative of $2\tilde{F}_{0}$ with respect to the frequency, and the terms $\tilde{F}_{4,0},\tilde{F}_{4,1}$ are affected by the correction terms $V^{\rm\pm}_{\rm 2nd},V^{\rm-+}_{\rm 2nd}$ and $V^{\rm s\pm}_{\rm 1st}$ in Eq. (46). The QNM correction $\delta\omega_{\rm grav}$ has two branches

### v2

@@ Abstract
We investigate black holes and gravitational perturbations when both the scalar Gauss-Bonnet and dynamical Chern-Simons gravity sectors coexist in addition to the Einstein-Hilbert term, and both sectors are coupled to a single canonically normalized scalar field. The presence of the scalar Gauss-Bonnet gravity sector allows the scalar field to possess a non-vanishing background solution, resulting in additional couplings between odd- and even-type gravitational perturbations arising from the dynamical Chern-Simons gravity sector. We illustrate the impact of these even-odd gravitational couplings in gravitational perturbations around a static spherically symmetric black hole. Although the couplings between the odd and even-type gravitational perturbations are known to appear in purely tensorial gravity theories with higher-curvature corrections, we demonstrate it in scalar-tensor theories.
@@ I Introduction
General relativity (GR) faces several challenges of the late-time acceleration Riess et al. (1998); Perlmutter et al. (1999), the formation of black hole singularities Penrose (1965); Hawking and Penrose (1970); Hawking (1976), and the quantization of gravity at both low- and high-energy scales. These strongly suggest that the theory of gravity should be modified from general relativity. In most cases, the modification of gravity involves adding extra degrees of freedom to the metric. In fact, the simplest modification is to introduce a non-minimal coupling of the metric to a scalar field, i.e., a scalar-tensor theory. As a typical example, we focus both on the scalar Gauss-Bonnet (sGB) gravity Metsaev and Tseytlin (1987); Kanti et al. (1996); Green and Schwarz (1984); Antoniadis et al. (1992) and the dynamical Chern-Simons (dCS) gravity Lue et al. (1999); Jackiw and Pi (2003); Alexander and Yunes (2009). Both theories of gravity are known as low-energy theories of string theory Mignemi and Stewart (1993); Antoniadis et al. (1994); Mignemi (1995); Kanti et al. (1996); Daniel and Jenks (2024). The sGB gravity sector breaks the four-dimensional topological invariance of the Gauss-Bonnet term by coupling the dynamical scalar field, whereas the dCS gravity sector has a parity-violating coupling between the scalar field and the Pontryagin term. [^1] We note that the sGB gravity is a healthy classical theory included in the Horndeski theory which is the most general theory with the second-order equations of motion Horndeski (1974); Deffayet et al. (2011); Kobayashi et al. (2011). The dCS gravity should be regarded as a low-energy effective field theory because of the presence of ghost modes around a static spherically symmetric black hole Motohashi and Suyama (2012) and any spacetime Crisostomi et al. (2018). The black hole solutions and QNMs have been studied in the sGB gravity Yunes and Stein (2011); Blázquez-Salcedo et al. (2016, 2017, 2019); Konoplya et al. (2019); Zinhailo (2019); Churilova and Stuchlik (2020); Blázquez-Salcedo et al. (2020); Pierini and Gualtieri (2021, 2022) and the dCS gravity Yunes and Sopuerta (2008); Cardoso and Gualtieri (2009); Molina et al. (2010); Yunes and Stein (2011); Srivastava et al. (2021); Wagle et al. (2022); Blázquez-Salcedo et al. (2023), respectively, where the deviations from the GR case can be expressed as a series of small couplings. Recently, in the context of the effective field theory (EFT), corrections to black hole solutions and QNMs with small EFT couplings have been studied Endlich et al. (2017); Cardoso et al. (2018); Cano and Ruipérez (2019); de Rham et al. (2020); Nomura and Yoshida (2022); Cano et al. (2022a); Silva et al. (2023); Cano et al. (2023a); Cayuso et al. (2023); Cano et al. (2023b); Melville (2024).
@@
In this paper, we study black holes and gravitational perturbations when both the sGB and dCS gravity sectors coexist, and both sectors are coupled to a single scalar field. The presence of the sGB gravity sector allows the scalar field to have a background value, leading to additional new couplings between the odd and even-type gravitational perturbations appearing through the dCS gravity sector. We illustrate the impact of the even-odd gravitational couplings in the perturbations around the static spherically symmetric black hole solutions. While the couplings between the odd- and even-type gravitational perturbations are known to appear in purely tensorial gravity theories with higher-curvature corrections Cardoso et al. (2018); de Rham et al. (2020); Cano et al. (2022b, 2023a), we demonstrate it in scalar-tensor theories.
@@ II.2 Static spherically symmetric black hole solutions
The thermodynamics are studied in the context of black holes in the Horndeski theory Minamitsuji and Maeda (2023).555 For fixed $\epsilon,b_{1}$ and $M$, the black hole solution in our setup is uniquely determined. This implies that the solution is thermodynamically stable, which is consistent with the results in Sec. IV where the quasinormal modes correspond to damped oscillations.
@@ III Gravitational Perturbations
In this section, we study gravitational perturbations around the static spherically symmetric background solutions $\bar{g}_{\mu\nu}$ and $\bar{\phi}$ in Eqs. (4) and (5) with Eqs. (7)–(14). The perturbed scalar field is given by
@@
where $h_{0},~{}h_{1},H_{0},~{}H_{1},~{}H_{2}$, and $K$ are functions of $r$. The odd (--) mode components $h_{0},~{}h_{1}$ and the even ($+$) mode components $H_{0},~{}H_{1},~{}H_{2},K$ are not independent, but related through the equations of motion. In our setup, the metric perturbations have two helicity-2 modes.
@@ III.1 Master equations
The explicit forms of ${\bm{\Psi}}$, $\Delta{\bm{c}}^{2}$, and ${\bm{V}}_{\rm correction}$ are written in App A. Since $\Delta{\bm{c}}^{2}$ vanishes at the horizon, the propagation speed of each mode coincides with the speed of light. This implies that information inside the horizon cannot propagate outside of the horizon. The non-diagonal components of the potential matrix in Eq. (31) represent the mode mixing. We note that the scalar-odd and scalar-even coupling terms at ${\cal O}(\epsilon^{2})$ in Eq. (31) vanish because the next leading terms appear at ${\cal O}(\epsilon^{3})$. Considering both sGB and dCS couplings, new terms, $V^{-+}_{\rm 2nd}$, can appear in the master equations. In particular, the even-odd gravitational couplings have been known to be a generic feature of purely tensorial gravity theories with higher-curvature corrections Cardoso et al. (2018); de Rham et al. (2020); Cano et al. (2022b, 2023a). We find that this is characteristic not only in purely tensorial gravity theories but also when both the sGB term and the dCS term are present. In the context of cosmology, it has been known that the even-odd gravitational couplings can appear in a similar system where both sGB and dCS gravity sectors coexist Satoh et al. (2008).
@@ IV Quasinormal Modes
QNMs are the characteristic oscillation modes of perturbations around black holes. The quasinormal frequencies are defined by the solutions of $\omega$ in the master equations imposing the purely ingoing condition at the horizon and the purely outgoing condition at infinity. [^7] In the context of initial value problems, QNMs are defined as poles in Green’s function and they correspond to the purely ingoing condition at the horizon and the purely outgoing condition at infinity for massless fields in asymptotically flat black holes Leaver (1986); Nollert and Schmidt (1992); Andersson (1995, 1997); Kokkotas and Schmidt (1999); Nollert (1999); Berti et al. (2009); Konoplya and Zhidenko (2011). The Leaver method Leaver (1985) is used to calculate the quasinormal frequencies. This calculation method is briefly presented for both decoupled and coupled systems. In the decoupled system, the master equation is described by a single master variable $\Psi$,
@@
where ${\rm d}r/{\rm d}r_{\ast}=f$. The wave function $\Psi$ of this form satisfies the QNM boundary condition at the horizon thanks to the prefactor $e^{-i\omega r_{\ast}}f^{-2i\omega r_{\rm H}}$. Substituting this ansatz into the master equation (34) yields the recursion relation for the coefficients $a_{j}$. For a given $a_{0}$, the higher-order coefficients $a_{j}~{}(j=1,~{}2,~{}\cdots)$ can be determined by the recursion relation888 Since the boundary condition at the horizon has already been imposed, the first term $a_{0}$ determines the solution of the master equation. in the form $a_{j}=G_{j}(\omega)a_{0}$, where $G_{j}(\omega)$ is a function in $j$ and $\omega$. If the wave function (33) satisfies the QNM boundary condition at infinity $f=1$, then the sum of the series $\sum^{\infty}_{j=0}a_{j}$ should converge. Such a solution can be derived approximately by imposing $a_{j}=0$ for sufficiently large $j$. Thus, we need to solve $G_{j}(\omega)=0$ for sufficiently large $j$, and the quasinormal frequencies are given by those solutions Leaver (1985). In GR, for $\ell=2$ and the fundamental modes, the quasinormal frequency for the scalar field is $(0.96728774-0.19359239i)/(2M)$ Berti et al. (2009) where $M$ is the black hole mass, and for the odd and even gravitational modes, it is $(0.74734337-0.17792463i)/(2M)$ Chandrasekhar and Detweiler (1975); Berti et al. (2009). In the case of non-GR gravity theories, if the master equation is expressed as a perturbation from the GR case, the quasinormal frequencies can be obtained as small deviations from the GR values Cardoso et al. (2018, 2019); McManus et al. (2019); Cano and Ruipérez (2019); de Rham et al. (2020); Kimura (2020); Nomura and Yoshida (2022); Cano et al. (2022a); Franchini and Völkel (2023); Cano et al. (2023a); Cayuso et al. (2023); Cano et al. (2023b); Hatsuda and Kimura (2024).
In a coupled system, the master equations have non-diagonal parts of the potential matrix999 Changing the master variable, the master equations (27) can be written in the form of Eq. (34) (see App. A in Cardoso et al. (2019)).
@@
where $\Omega^{i}_{\rm Sch}~{}(i={\rm s,~{}grav.})$ are the dimensionless Schwarzschild quasinormal frequencies normalized by the Schwarzschild radius for the scalar field, the odd and even gravitational modes.101010 Practically, we first calculate $\delta\bar{\omega}_{i}$ with $\omega=\Omega_{\rm Sch}^{i}/r_{\rm H}+\epsilon^{2}\delta\bar{\omega}_{i}$, and then we can obtain $\delta\omega_{i}$ in Eq. (37) as $\delta\omega_{i}=\delta\bar{\omega}_{i}-\delta R_{\rm H}\delta\Omega_{\rm Sch}^{i}/(2M)$, where $r_{\rm H}=2M(1+\epsilon^{2}\delta R_{\rm H})$. For the fundamental modes, the values of $\Omega^{i}_{\rm Sch}$ are given by $\Omega^{\rm s}_{\rm Sch}=0.96728774-0.19359239i$ and $\Omega^{\rm grav}_{\rm Sch}=0.74734337-0.17792463i$. Substituting these ansatzes into Eq. (36) in our system and expanding them with respect to $\epsilon$, we obtain the equations determining $\delta\omega_{i}$.
@@
where $F_{0},F_{2,0}$, and $F_{2,1}$ are functions of $\Omega^{\rm s}_{\rm Sch}$. The lowest-order value $\Omega^{\rm s}_{\rm Sch}$ is determined by the equation $F_{0}=0$. The second order ${\cal O}(\epsilon^{2})$ expression of Eq. (38) determines the value of $\delta\omega_{\rm s}$. We note that the term $F_{2,1}$ is the derivative of $F_{0}$ with respect to the frequency, and the term $F_{2,0}$ is affected by the correction terms $V^{\rm s}_{\rm 2nd}$ and $V^{\rm s\pm}_{\rm 1st}$ in Eq. (31). The QNM correction $\delta\omega_{\rm s}$ is given by
@@
The corrections to the lowest-order values of the odd and even gravitational modes are determined by the self-couplings of those modes $V^{\rm-}_{\rm 2nd}$ and $V^{\rm+}_{\rm 2nd}$, and the even-odd gravitation coupling $V^{\rm-+}_{\rm 2nd}$. Equation (36) is given by
@@
The terms $\tilde{F}_{0},\tilde{F}_{2,0},\tilde{F}_{2,1},\tilde{F}_{4,0},\tilde{F}_{4,1},$ and $\tilde{F}_{4,2}$ are functions of $\Omega^{\rm grav}_{\rm Sch}$. We solve the above equation order by order in $\epsilon$. The lowest-order quasinormal frequency of the gravitational modes $\Omega^{\rm grav}_{\rm Sch}$ is determined by $\tilde{F}_{0}=0$. Since the spectra of even and odd gravitational modes are degenerate at the lowest order, the terms $\tilde{F}_{2,0}$ and $F_{2,1}$, which are proportional to the first derivative of $\tilde{F}_{0}$, also vanish. Thus, the QNM correction $\delta\omega_{\rm grav}$ is determined from Eq. (40) with ${\cal O}(\epsilon^{4})$. We note that the term $\tilde{F}_{4,2}$ is the second derivative of $2\tilde{F}_{0}$ with respect to the frequency, and the terms $\tilde{F}_{4,0},\tilde{F}_{4,1}$ are affected by the correction terms $V^{\rm\pm}_{\rm 2nd},V^{\rm-+}_{\rm 2nd}$, and $V^{\rm s\pm}_{\rm 1st}$ in Eq. (31). The QNM correction $\delta\omega_{\rm grav}$ has two branches